OpenStudy (anonymous):

you have a part time job. you work for 3 hours on friday and 6 hours on saturday. you also receive an allowance of $20 per week. you earn $92 per week how much doyou earn per hour at your part time job

OpenStudy (whpalmer4):

income = allowance + job hours * job pay rate $92 = $20 + (3+6)*pay rate $92 = $20 + 9 * pay rate $72 = 9 * pay rate $8 = pay rate